# Data for Mathematical Copilots: Better Ways of Presenting Proofs for Machine Learning

### v1

@@ 3.2 Distinct Issues
A natural way to circumvent the difficulty of evaluating natural language proofs is by making use of formal language datasets. Instead of having an LLM generate natural language mathematical proofs, one can train it to generate proofs in a verifiable formal language such as Lean or Isabelle.
Therefore, existing libraries of formal proofs can play a crucial role in the development of machine learning-based automatic theorem provers. Sections 3.2.2 and 3.2.3 will provide details on how some of the largest existing collections of formal proofs have been and are being created. This is followed by an analysis of how this process affects the use of formal proof libraries as evaluation datasets; with some remarks relevant to using formal libraries for training.
More concretely, Isabelle’s Archive of Formal Proofs (AFP) and Lean’s Mathlib will be investigated. This choice is due to the authors’ familiarity with these libraries but many of the points mentioned below will also be valid (at least to some extent) for other systems.
One particular issue relevant to all datasets is their scale. The lack of a large-scale dataset consisting of pairs of formal and informal mathematics is a significant bottleneck for autoformalization. Current datasets, such as miniF2F are on the order of a few hundred datapoints, which are enough for assessing autoformalization, but not for training models to support autoformalization. A recent effort (Ying et al., 2024) scales this to 57k datapoints of pairs. Nonetheless, this scale is not yet comparable to the scale at which LLMs can be trained on informal mathematics.
@@ 3.2.1 Lack of Trusted Automatic Evaluation
No large machine learning dataset on mathematics currently exists that allows an arbitrary proof of an AI system to be checked for correctness against the gold-standard proof in the evaluation dataset. A proof-of-concept, highlighted by the IMO Small Challenge888www.imo-small-challenge.io, shows how partial automation might be achieved, where a necessary test for correctness is carried out using detailed proof annotation. This builds on the idea that a system like an LLM, combined with a deterministic form of matching, can at least exclude incorrect proof candidates. If the gold truth is sufficiently rich, the annotations help the LLM break down a proof and assist it in understanding its main features (accepting that subtler points of proofs may still be currently out of reach for LLM understanding).
Some success has been attained in using an LLM to grade (or teach) another LLM (Eldan and Li, 2023; Mukherjee et al., 2023). Yet, for mathematics, we believe that this is not necessarily the right approach forward yet. The current generation of models do not have a sufficiently high performance on mathematics to be used as graders: for tasks that would often be deemed simpler than mathematics, this approach can fail (Wang et al., 2023), although for solving programming puzzles this approach was successful (Haluptzok et al., 2023). Anecdotal evidence suggests that adapting the approach to proofs is challenging because of the multitude of potential proof variations for a single statement, which gives rise to a diverse set of textual expressions.
At present, grading can, at best, be automated by the use of detailed human annotation for each problem in combination with LLM assistance based on such annotation.
@@ 3.2.2 Tool Misalignment
One cause of this discrepancy is that side conditions deemed trivial in conventional mathematical literature have to be formalized as diligently as the rest of the mathematical text. On the other hand, proving many statements considered routine by mathematicians can be mechanized.
For example, proving that a function is continuous often corresponds to choosing the correct subset of a few relevant lemmas and this can be achieved with a simple search algorithm, such as Lean’s continuity tactic.
Next to this domain-specific formalization aid, there are also general-purpose automation tools that can be used to formalize certain simple statements. For example, Isabelle’s Sledgehammer tool (Paulson and Susanto, 2007; Meng and Paulson, 2008) translates the given statement so that it can be understood by a number of external automatic theorem provers (ATPs). The ATPs, in turn, attempt to prove the statement and send the proof back to Isabelle.
@@
The impact of these mechanisms needs to be kept in mind when formal datasets such as Lean’s Mathlib or Isabelle’s AFP are used for LLM evaluation. In particular, it is likely that current datasets overly reward models for their ability to deploy the language-specific automation, rather than complete “understanding” of the underlying mathematics. As some evidence of this, Hu et al. (2024) show that LLM performance drops significantly when common automation tactics are turned off, even for theorems that have human-written proofs that do not rely on the automated tactics.
@@
A natural way to turn the AFP into a proof dataset is by splitting up all formal proofs into their individual steps (Jiang et al., 2021). However, by the above characterization, such a dataset might be skewed towards being tractable for automation because of the particular way in which AFP theorems have come into existence. This can make it hard to assess the level of difficulty in the problem set.
@@
We would like to stress that we believe that the use of formal mathematics datasets such as the AFP is a valid evaluation tool. The results of recent publications  (First et al., 2023; Jiang et al., 2022; Mikuła et al., 2024) are especially promising, and we are looking forward to their continuation. Nevertheless, we believe that future evaluation should carefully consider the impact of existing automation in ITPs on their libraries.
@@ 3.2.3 Non-Trivial Data Duplication
Following the discussion above, we recommend developing methodologies that ensure that train/test splits adequately measure a model’s ability to generalize. One direction is to test on new formalizations that occur after a model has been trained, as was explored in Hu et al. (2024). Nevertheless, care should be taken to ensure that the mathematics being formalized does not occur in previous projects. Hence, we particularly recommend evaluating on domains that have not previously been formalized.
@@ 4 The Future – Novel Datasets to Support Mathematical Copilots
In this section, we cover various aspects of mathematical research practice that, to date, are not covered at all by any of the existing datasets. This contrasts with the previous section on mathematical datasets that covered parts of mathematical practice, albeit with the noted limitations. Current datasets exclusively are focussed on publishing the results rather than the intermediate steps that mathematician goes through in the process of devising a proof. The different types of workflows, the proof-counterexample dialectic, and proof transfer,
etc., are all examples of intermediate processes that do not have clear representations in data that can be used to fine-tune LLMs. Contrasting with the previous section, where we made specific recommendations on how to improve existing datasets, in this section, we do not make specific recommendations, as our recommendation is to simply devise such datasets. While a few inroads have been made into non-proof-based datasets, see Section 3.1.3 these are still very much at their inception, and, to our knowledge, the GHOSTS dataset is the only attempt to date to try to “data-ify” some aspects (solely related to mathematical literature search) of daily mathematical process.
@@ 4.1 Mapping Mathematical Workflows to Data
While a workflow (indicated with arrows below) can, in principle, easily be mapped to data by converting it to $n$ tuples of datapoints (where $n$ denotes the number of workflow steps) and filling text in between, in practice, in certain cases, issues may arise. We give examples of both global and local workflows below and highlight both examples where these can be transformed to data in a straightforward manner, as well as cases in which this transformation is more difficult. We urge the community to explore such translation of workflows to data further.
@@
We list below examples from two subfields of mathematics (differential equations and knot theory/low-dimensional topology).
  • Suppose, for instance, one has a first-order ODE $\partial_{t}u=F(u)$ with some initial data $u(0)=u_{0}$ and wants to know how the solution grows in time. A standard technique is to introduce a key functional $E(u)$ of the solution (often something like an “energy” or “Lyapunov functional”) and then compute the derivative $\partial_{t}E(u)$ using the ODE and the chain rule (for PDEs, one often has to also perform several times integration by parts). Then, one bounds this derivative as best one can.
    The goal is to reach some differential inequality of Gronwall type, e.g., $\partial_{t}E(u)\leq CE(u)$, so that a Gronwall-type lemma may be applied (but it is not always precisely the Gronwall lemma, but something similar).
@@
    We note that individual steps of this may be solved using symbolic-numeric methods, which opens the possibility of using a tool-integrated-reasoning (TIR) approach to enable an LLM to discharge these proof steps to symbolic or numeric tools to, e.g., compute derivatives symbolically. Thus, one option to encapsulate these workflow steps is in datapoints whose core form is
@@
    and where there is filler text between the raw mathematical objects (i.e., the ODE, the functional, etc.). The raw mathematical objects should be symbolically or numerically generated so that custom filler text can be added in between them, and an LLM can observe a range of ways in which to speak and textually connect a fixed tuple of mathematical objects.
@@
We now consider typical workflows in low-dimensional topology, which is an area of mathematics that has a unique flavor and uses tools from geometry, algebra, PDEs, group theory, combinatorics, and mathematical physics. Hence, some of these workflows are more challenging to convert to data. Low-dimensional topology is the study of $n$-manifolds (certain topological spaces that are locally homeomorphic to $\mathbb{R}^{n}$) of dimension $n\leq 4$. A knot is a simple closed curve embedded in $\mathbb{R}^{3}$. Knot theory plays a fundamental role in low-dimensional topology as every 3- and 4-manifold can be represented by a framed link; i.e., a collection of knots labeled by integers, called a Kirby diagram (Kirby, 1978). We refer the reader to Juhász (2023) for more detail on low-dimensional topology and knot theory and explanations of the mathematical terms used below.
@@
    To show that two representations of the same knot or manifold are equivalent, one finds a sequence of certain moves connecting them, which can be Reidemeister moves (Reidemeister, 1927) in case of knot diagrams, Markov moves for braids, or Kirby moves in case of Kirby diagrams. These search problems lend themselves to techniques such as reinforcement learning; see (Gukov et al., 2021). The workflow steps can be encoded by a sequence of representations of the object, such that consecutive representations are related by one of the standard moves. Software packages such as SnapPy (Culler et al., 2024) can be used to check whether these moves are valid and can also list valid moves. We can hence represent the workflow as follows:
@@
  • To show two objects are inequivalent, one defines invariants, which are typically algebraic objects (numbers, polynomials, or groups) that are unchanged by the above moves. These moves often also capture important topological properties of these objects. The most classical invariants are homology groups and the fundamental group. The Alexander polynomial is a knot invariant derived from the fundamental group of the knot complement. Representations of the knot group give rise to the more modern twisted Alexander polynomials. A recent knot invariant rooted in representation theory is Khovanov homology (Khovanov, 2000). Floer homology (Ozsváth and Szabó, 2004) and gauge theory (Witten, 1994) give rise to highly sophisticated knot and 3- and 4-manifold invariants. These invariants often give lower bounds on hard-to-compute topological quantities such as the 3- or 4-genus of a knot. SnapPy is capable of computing many of these knot invariants when run in SageMath (The Sage Developers, 2024). A typical workflow would thus be as follows:

### v2

@@ 3.2 Distinct Issues
A natural way to circumvent the difficulty of evaluating natural language proofs is by making use of formal language datasets. Instead of having an LLM generate natural language mathematical proofs, one can train it to generate proofs in a verifiable formal language such as Lean or Isabelle. Therefore, existing libraries of formal proofs can play a crucial role in the development of machine learning-based automatic theorem provers. Sections 3.2.2 and 3.2.3 will provide details on how some of the largest existing collections of formal proofs have been and are being created. This is followed by an analysis of how this process affects the use of formal proof libraries as evaluation datasets; with some remarks relevant to using formal libraries for training. More concretely, Isabelle’s Archive of Formal Proofs (AFP) and Lean’s Mathlib will be investigated. This choice is due to the authors’ familiarity with these libraries but many of the points mentioned below will also be valid (at least to some extent) for other systems. One particular issue relevant to all datasets is their scale. The lack of a large-scale dataset consisting of pairs of formal and informal mathematics is a significant bottleneck for autoformalization. Current datasets, such as miniF2F are on the order of a few hundred datapoints, which are enough for assessing autoformalization, but not for training models to support autoformalization. A recent effort (ying2024lean) scales this to 57k datapoints of pairs. Nonetheless, this scale is not yet comparable to the scale at which LLMs can be trained on informal mathematics.
@@ 3.2.1 Lack of Trusted Automatic Evaluation
No large machine learning dataset on mathematics currently exists that allows an arbitrary proof of an AI system to be checked for correctness against the gold-standard proof in the evaluation dataset. The Open Proof Corpus collects expert judgments on model-generated natural-language proofs and reports large gaps between boxed-answer accuracy and proof validity (openproofcorpus2025). A proof-of-concept, highlighted by the IMO Small Challenge888www.imo-small-challenge.io, shows how partial automation might be achieved, where a necessary test for correctness is carried out using detailed proof annotation. This builds on the idea that a system like an LLM, combined with a deterministic form of matching, can at least exclude incorrect proof candidates. If the ground truth is sufficiently rich, the annotations help the LLM break down a proof and understand its main features (accepting that subtler points of proofs may still be currently out of reach of LLMs’ understanding).
Some success has been attained in using an LLM to grade (or teach) another LLM (eldan2023tinystories; mukherjee2023orca). Yet, for mathematics, we believe that this is not necessarily the right approach forward. The current generation of models does not have a sufficiently high performance on mathematics to be used as graders: for tasks that would often be deemed simpler than mathematics, this approach can fail (wang2023large), although for solving programming puzzles this approach was successful (haluptzok2022language). Anecdotal evidence suggests that adapting the approach to proofs is challenging because of the multitude of potential proof variations for a single statement, which gives rise to a diverse set of textual expressions. At present, grading can, at best, be automated by the use of detailed human annotation for each problem in combination with LLM assistance based on such annotation.
@@ 3.2.2 Tool Misalignment
One cause of this discrepancy is that side conditions deemed trivial in conventional mathematical literature have to be formalized as diligently as the rest of the mathematical text. On the other hand, proving many statements considered routine by mathematicians can be mechanized. For example, proving that a function is continuous often corresponds to choosing the correct subset of a few relevant lemmas and this can be achieved with a simple search algorithm, such as Lean’s continuity tactic. Next to this domain-specific formalization aid, there are also general-purpose automation tools that can be used to formalize certain simple statements. For example, Isabelle’s Sledgehammer tool (sledgehammer2007; sledgehammer2008) translates the given statement so that it can be understood by a number of external automatic theorem provers (ATPs). The ATPs, in turn, attempt to prove the statement and send the proof back to Isabelle.
@@
The impact of these mechanisms needs to be kept in mind when formal datasets such as Lean’s Mathlib or Isabelle’s AFP are used for LLM evaluation. In particular, it is likely that current datasets overly reward models for their ability to deploy the language-specific automation, rather than complete “understanding" of the underlying mathematics. As some evidence of this, hu2024minictxneuraltheoremproving show that LLM performance drops significantly when common automation tactics are turned off, even for theorems that have human-written proofs that do not rely on the automated tactics.
@@
A natural way to turn the AFP into a proof dataset is by splitting up all formal proofs into their individual steps (jiang2021lisa). However, by the above characterization, such a dataset might be skewed towards being tractable for automation because of the particular way in which AFP theorems have come into existence. This can make it hard to assess the level of difficulty in the problem set.
@@
We would like to stress that we believe that the use of formal mathematics datasets such as the AFP is a valid evaluation tool. The results of recent publications (first2023baldur; thor; magnushammer) are especially promising, and we are looking forward to their continuation. Nevertheless, we believe that future evaluation should carefully consider the impact of existing automation in ITPs on their libraries.
@@ 3.2.3 Non-Trivial Data Duplication
Following the discussion above, we recommend developing methodologies that ensure that train/test splits adequately measure a model’s ability to generalize. One direction is to test on new formalizations that occur after a model has been trained, as was explored in hu2024minictxneuraltheoremproving. Nevertheless, care should be taken to ensure that the mathematics being formalized does not occur in previous projects. Hence, we particularly recommend evaluating on domains that have not previously been formalized.
@@
3.2.4 Version Fragility and Tactic Stability
Formal proof corpora are often treated as static datasets consisting of a theorem statement plus a proof script. In practice, they are tightly coupled to a moving toolchain (involving a specific version of the prover, a particular commit of the library, and details of automation implied by this combination). This can make them fairly brittle as datasets: even small library changes can invalidate proofs or silently change what intermediate tactics do.
A common failure mode is “tactic instability”, where a proof uses a flexible tactic (in Lean, most notably simp) and then applies a subsequent tactic whose success depends on the exact transformed goal. Because the set of parameters used by the automated tactic (e.g., lemmas used by simp) evolves, the transformed goal may drift over time, making the downstream step brittle even if the overall theorem statement is unchanged.
Recommendations.
Recent work in Mathlib [^11] See https://leanprover-community.github.io/mathlib4_docs/Mathlib/Tactic/Linter/FlexibleLinter.html for details. introduces a stability linter (“FlexibleLinter”) that flags precisely these patterns, encouraging authors to make tactic behavior more explicit (e.g., by specifying a minimal set of lemmas used in the simplification). As a dataset design principle, proofs that pass such stability linting are better candidates for long-lived benchmarks, since their execution is less sensitive to future library drift.
3.2.5 Hidden Context Dependence: Local Instances and Attribute State
Even in a fully formal setting, a dataset item is seldom determined by the goal alone. In Lean, imports, local instances, and attribute state can change the meaning of a line of proof text. This matters for dataset construction because “the same” proof script can behave differently when replayed under a different environment.
Example (instance “pollution”).
In measure-theoretic developments in Lean, the proposition MeasurableSet s is parameterized by an implicit measurable-space instance. If a proof introduces a local measurable space m0 but an ambient instance mW remains in scope, MeasurableSet may silently refer to mW unless the instance is pinned:
      -- Fails: ‘MeasurableSet‘ uses the ambient instance, not the local one    have h_bad : MeasurableSet s := by     simpa using hs        -- Works: pin the intended instance explicitly    have h_ok : @MeasurableSet alpha m0 s := by     simpa [m0] using hs
This illustrate that storing only the final proof term (or only the local goal) in a dataset can be insufficient to replay a proof deterministically, because replay depends on the surrounding instance environment and imported modules. For more examples from various areas of math, see the reference patterns contained in the Claude Code skill for Lean 4 (FreerLean4Skills).
Example (context-dependent simplification).
Automation can also be context-sensitive in a more mundane way. Automated tactics such as an instance of simp in Lean can succeed or fail depending on hypotheses introduced earlier in the proof:
      -- Works with extra context in scope    filter_upwards [hw, hu_eps, hv_eps] with w hu hv    simp [Real.norm_eq_abs]        -- May fail without that context, even if the syntactic goal looks similar    simp [Real.norm_eq_abs]
This suggests that proof snippets that are extracted without their surrounding context (imports, local hypotheses, attributes) can be underspecified as dataset items for training or evaluation.
3.2.6 Resource Sensitivity: Heartbeats and Elaboration as Part of the Task
In interactive theorem provers, “correctness” is not clearly delineated from performance. A proof that compiles only under extremely permissive resource limits (such as “heartbeats”, recursion depth, or wall-clock time) is often not usable in real formalization workflows, and may not be reproducible across environments (especially if the high resource bounds are not clearly specified).
Lean exposes heartbeat and other resource limits directly:
      set_option maxHeartbeats 40000    set_option synthInstance.maxHeartbeats 40000    set_option maxRecDepth 2000
These resource bounds can be exceeded by complex elaborations or typeclass searches. Sometimes this is due to the complexity of the underlying mathematics, but other times it is from “fighting the elaborator”. For instance, leaving instance arguments implicit can trigger expensive unification/typeclass search, sometimes consuming hundreds of thousands of heartbeats before timing out. This suggests that datasets and benchmarks should treat resource budgets as first-class metadata, and report resource usage alongside whether the code passed or failed.
3.2.7 Redundant Scaffolding and Accidental Complexity in Proof Scripts
Formal corpora often contain a lot of proof-script structure that is not mathematically essential, including redundant tactic steps, arguments to automation that are not used, and lemma variants that exist only to satisfy local engineering constraints. As dataset artifacts, these patterns can dominate the training signal while teaching models brittle or unidiomatic behavior.
Example (no-op tactic after automation).
A common failure mode when replaying mined proof scripts is that one tactic already closes the goal, making subsequent steps ill-typed. For example, in Lean, this sequence can fail:
      simp -- closes the goal    aesop -- error: no goals to be solved
Example (lintable redundancy in automation arguments).
Even when scripts compile, they often include extra automation arguments that are unused in practice. Mathlib-style linting can flag such redundancy and enable construction of a cleaned corpus:
      -- Linter: unused simp arguments (suggest removing to reduce clutter)    simp [pow_two, mul_assoc, mul_add, mul_comm, add_comm, add_left_comm]
In light of this, it may be useful for datasets to include both a raw corpus (for realism) and a linter-cleaned corpus (for learning canonical structure).
@@ 4 The Future – Novel Datasets to Support Mathematical Copilots
In this section, we cover various aspects of mathematical research practice that, to date, are not covered at all by any of the existing datasets. This contrasts with the previous section on mathematical datasets that covered parts of mathematical practice, albeit with the noted limitations. Current datasets are exclusively focused on publishing the results rather than the intermediate steps that a mathematician goes through in the process of devising a proof. The different types of workflows, the proof-counterexample dialectic, and proof transfer, etc., are all examples of intermediate processes that do not have clear representations in data that can be used to fine-tune LLMs. Contrasting with the previous section, where we made specific recommendations on how to improve existing datasets, in this section, we do not make specific recommendations, as our recommendation is to simply devise such datasets. While a few inroads have been made into non-proof-based datasets, see Section 3.1.3 these are still very much at their inception, and, to our knowledge, the GHOSTS dataset is the only attempt to date to try to “data-ify” some aspects (solely related to mathematical literature search) of daily mathematical process.
An example of proof transfer is given by the Ax–Grothendieck theorem, which states that every injective polynomial map $\mathbb{C}^{n}\to\mathbb{C}^{n}$ is surjective; rather than proving this directly over $\mathbb{C}$, one observes that the statement is trivial for finite fields – any injective map on a finite set is bijective – and then uses model-theoretic compactness to transfer the result to algebraically closed fields of characteristic zero. Here, one does not redo the argument in the new structure directly; instead, one uses model theory to move the truth of a sentence from one place to another.
@@ 4.1 Mapping Mathematical Workflows to Data
Simple methods as precursors.
Even relatively simple machine learning approaches have produced genuine mathematical discoveries. Wagner (Wagner:2021Constructing) demonstrated that reinforcement learning, framed as a single-player graph-building game, can find counterexamples to open conjectures in extremal combinatorics. PatternBoost (Charton:2024Patternboost), built on Karpathy’s character-level language model makemore (karpathy2022makemore), alternates between local search and transformer-based generation to discover constructions that refuted a 30-year-old conjecture. These methods succeed precisely because they operate in domains where candidate solutions can be automatically evaluated. This is a form of verifiable feedback that sidesteps the hallucination problems plaguing open-ended generation. They are precursors to more general workflows that combine neural generation with symbolic verification and human mathematical insight.
@@
While a workflow (indicated with arrows below) can, in principle, easily be mapped to data by converting it to $n$ tuples of datapoints (where $n$ denotes the number of workflow steps) and filling text in between, in practice, in certain cases, issues may arise. We give examples of both global and local workflows below and highlight both examples where these can be transformed to data in a straightforward manner, as well as cases in which this transformation is more difficult. We urge the community to explore such translation of workflows to data further. Future work should address how such workflow datasets might be structured, how symbolic or numeric methods can be integrated into individual workflow steps, and which parts of these workflows are best left to humans versus delegated to a mathematical copilot. Once such datasets exist, LLMs can be trained on them using standard methods such as expert iteration (DBLP:conf/nips/AnthonyTB17; havrilla2024teachinglargelanguagemodels).
Inference-time compute in reinforcement learning needs problem analyses.
Several recent math reasoning systems use substantial inference-time compute, for example, by sampling many candidate solutions, running search over intermediate steps, or iterating tool calls, and then using reinforcement learning to train policies that decide which branches to expand. In domains with automatic verification, these loops can rely on a binary reward signal. For most informal mathematical tasks, that signal is too sparse: a model receives feedback only after a long chain of steps, and failure provides little information about where the attempt went off track.
One remedy is to include, alongside each problem statement, a structured problem analysis that makes intermediate progress measurable. Concretely, this can include a normalized restatement of the problem, the relevant definitions and constraints, a decomposition into subgoals that a solver would aim to establish, and examples of admissible intermediate claims (or failure modes) that an evaluator can recognize. These annotations can be authored by experts or drafted with model assistance and then checked, but they are rarely present in current math benchmarks. Without this layer, RL methods that spend compute at inference time are pushed toward brute-force sampling on a sparse objective, which does not scale to longer proofs or to tasks whose correctness is not cheaply checkable.
@@
We list below examples from several distinct subfields of mathematics (differential equations, knot theory/low-dimensional topology, and number theory).
  • Suppose, for instance, one has a first-order ODE $\partial_{t}u=F(u)$ with some initial data $u(0)=u_{0}$ and wants to know how the solution grows in time. A standard technique is to introduce a key functional $E(u)$ of the solution (often something like an “energy” or “Lyapunov functional”) and then compute the derivative $\partial_{t}E(u)$ using the ODE and the chain rule (for PDEs, one often has to also perform several times integration by parts). Then, one bounds this derivative as best one can. The goal is to reach some differential inequality of Gronwall type, e.g., $\partial_{t}E(u)\leq CE(u)$, so that a Gronwall-type lemma may be applied (but it is not always precisely the Gronwall lemma, but something similar).
@@
    We note that individual steps of this may be solved using symbolic-numeric methods, which opens the possibility of using a tool-integrated-reasoning (TIR) approach to enable an LLM to discharge these proof steps to symbolic or numeric tools to, e.g., compute derivatives symbolically. Thus, one option to encapsulate these workflow steps in data is by using datapoints whose form is
@@
    with potentially filler text between the raw mathematical objects contained in the datapoint (i.e., the ODE, the functional, etc.). The raw mathematical objects should be symbolically or numerically generated so that custom filler text can be added in between them, and an LLM can observe a range of ways in which to speak and textually connect a fixed tuple of mathematical objects.
@@
  • A commonly encountered workflow for evolutionary PDEs (i.e., PDEs whose solution evolves with a time variable $t$), analogous to the first-order ODE example given above, can be illustrated as follows. One wants to find a functional which will facilitate providing useful information about solutions of the PDE. To illustrate how such functionals are found in practice, we begin by considering the simple example of the linear wave equation $\partial_{tt}u-\Delta u=0$ on a bounded domain $\Omega\subset\mathbb{R}^{n}$, with smooth boundary $\partial\Omega$, subject to homogeneous Dirichlet boundary conditions $u=0$ on $\partial\Omega$ and initial conditions $u(x,0)=u_{0}(x)$, $\partial_{t}u(x,0)=u_{1}(x)$. (One can also study this problem on an unbounded domain $\Omega$ such as, for instance, $\mathbb{R}^{n}$, and impose that solutions must decay in the unbounded directions of $\Omega$.) To show energy conservation, we define the energy functional
    $E(t)=\frac{1}{2}\int_{\Omega}\Bigl[(\partial_{t}u)^{2}+|\nabla u|^{2}\Bigr]\,dx,$ (2)
    which represents the sum of kinetic energy and potential energy. Differentiating $E(t)$ in time under the integral sign, we obtain
    $\frac{dE}{dt}=\int_{\Omega}(\partial_{t}u)\,(\partial_{tt}u)\,dx+\int_{\Omega}\nabla u\cdot\nabla(\partial_{t}u)\,dx.$
    An integration by parts shifts the spatial derivative in the second term, while the boundary contribution vanishes because $u$ and hence $\partial_{t}u$ are zero on $\partial\Omega$. By substituting the wave equation $\partial_{tt}u=\Delta u$, the expression collapses to zero, showing that $\tfrac{dE}{dt}=0$. Thus
    $E(t)=E(0)\text{ for all }t>0.$ (3)
    More generally, on any finite time interval $[0,T]$, one can often prove an energy bound of the form $E(t)\leq C(T)\,E(0)$ for $0\leq t\leq T$, with $C(T)$ possibly being a constant function. When $C(T)=C>0$ is constant, we write
    $E(t)\leq CE(0)\text{ for all }t>0.$ (4)
    An inequality like (4) is sometimes referred to as an energy estimate. Notice that the conservative situation (3) satisfies the inequality (4) with $C=1$. The energy bound (4) can be proved for certain wave equations with non-constant coefficients and for some non-linearities (often under additional structural or smallness assumptions); it provides useful control on the solution. For example, it can be used to prove the uniqueness of solutions of the wave equation (evans2010pde, Section 2.4, Theorem 5), as well as finite propagation speed for solutions of the (constant-coefficient) wave equation (evans2010pde, Section 2.4, Theorem 6), and more generally under standard hyperbolicity assumptions.
    One can invert the previous perspective on this problem: if one had no prior knowledge of this equation and its properties, how could equations (3) or (4) be derived in the first place? In particular, to obtain an energy identity (and hence an energy estimate) for this model problem, one takes the equation $\partial_{tt}u-\Delta u=0$ and multiplies both sides of it by $\partial_{t}u$, and then integrates by parts on $0,T]\times\Omega$, where $T>0$. (The exact choice of this multiplier depends on the equation at hand; as we will see in the next paragraph, $u$ turns out to be a good multiplier for the heat equation. Moreover, multiple good choices of multipliers are usually known to exist for each equation, with different multipliers usually leading to different and useful inequalities, among which (4) can be viewed as one of the simplest examples.) Assuming that either $u=0$ on $\partial\Omega$ (Dirichlet, hence $\partial_{t}u=0$ on $\partial\Omega$) or $\partial_{\nu}u=0$ on $\partial\Omega$ (Neumann), after a few integrations by parts, one obtains exactly that the energy functional defined in (2) is conserved: $E(t)=E(0)$ for all $t\geq 0$. It turns out that for many more general wave equations of the form $Pu=F$ (where $P$ is a wave-type differential operator, e.g. $\partial_{tt}-\Delta$ with variable coefficients and lower-order terms, and $F$ is a forcing term), the same choice of multiplier $\partial_{t}u$ (meaning we multiply the equation by $\partial_{t}u$ and integrate by parts) yields an energy identity or inequality. Under additional structural assumptions, this can be turned into a uniform-in-time bound of the form (4). In practice, finding good multiplier functions and performing integration by parts in the resulting manipulations is one of the standard ways new inequalities are discovered.
    Let us modify the wave equation argument slightly to obtain an analogous result for the linear heat equation $\partial_{t}u-\Delta u=0$. In this case, as we will now see, one obtains
    $E(t)\leq E(0),\qquad t\geq 0$ (5)
    where $E(t)$ is defined to be $E(t)=\frac{1}{2}\int_{\Omega}u^{2}\,dx$; notice that this functional $E$ differs from the functional we used for the wave equation. As in the wave equation case, the inequality (5) can be discovered by multiplying both sides of the heat equation by an appropriate function (here, $u$), and then integrating on the domain $0,T]\times\Omega$. After a few uses of integration by parts, if $u=0$ on $\partial\Omega$ (Dirichlet) or $\partial_{\nu}u=0$, then one obtains, for $0\leq t\leq T$,
    $E(t)+\int_{0}^{t}\|\nabla u(t,\cdot)\|_{L^{2}(\Omega)}^{2}\,dt=E(0).$
    Since $\int_{0}^{t}\|\nabla u(t,\cdot)\|_{L^{2}(\Omega)}^{2}\,dt\geq 0$, one concludes (5). The inequality (5) can be used to prove the uniqueness of solutions to the heat equation; one can find in (evans2010pde, Section 2.3, Theorem 10) a proof of uniqueness of solutions forward in time, and (evans2010pde, Section 2.3, Theorem 11) proves a backward uniqueness property. (Unlike the wave equation, the notions of uniqueness forward and backwards in time do not “coincide” immediately for the heat equation.) The idea of manipulating a useful functional such as $E(t)$ goes by the name of “energy methods” in the PDE literature. The “energy method” for the heat equation serves as a template in the following sense: many more difficult non-linear PDEs (of which the Navier–Stokes system is one example) admit a version of this “multiply by the unknown and integrate” trick, and understanding it in the heat equation model helps one to understand those.
    The main insight in these “energy method” arguments for both equations was the introduction of the energy functional $E(t)$. We saw that for both the heat and wave equations, this functional can be derived by multiplying the equation with an appropriately chosen multiplier—specifically, $u$ for the heat equation and $\partial_{t}u$ for the wave equation—and then integrating by parts over the domain $0,T]\times\Omega$, with $T>0$. A mathematical copilot might successfully identify useful and new energy functionals by exploring various chosen multipliers and then attempting to derive useful quantities that can be brought under control by manipulating the resulting integrals. Newly discovered functionals which can be controlled may provide valuable new advances in mathematicians’ understanding of solutions to the corresponding equation(s).
@@
We now consider typical workflows in low-dimensional topology, which is an area of mathematics that has a unique flavor and uses tools from geometry, algebra, PDEs, group theory, combinatorics, and mathematical physics. Hence, some of these workflows are more challenging to convert to data. Low-dimensional topology is the study of $n$-manifolds (certain topological spaces that are locally homeomorphic to $\mathbb{R}^{n}$) of dimension $n\leq 4$. A knot is a simple closed curve embedded in $\mathbb{R}^{3}$. Knot theory plays a fundamental role in low-dimensional topology as every 3- and 4-manifold can be represented by a framed link; i.e., a collection of knots labeled by integers, called a Kirby diagram (Kirby:1978). We refer the reader to Juhasz2023book for more detail on low-dimensional topology and knot theory and explanations of the mathematical terms used below.
@@
    To show that two representations of the same knot or manifold are equivalent, one finds a sequence of certain moves connecting them, which can be Reidemeister moves (Reidemeister:1927Abhandlungen) in case of knot diagrams, Markov moves for braids, or Kirby moves in case of Kirby diagrams. These search problems lend themselves to techniques such as reinforcement learning; see (Gukov:2021Learning). The workflow steps can be encoded by a sequence of representations of the object, such that consecutive representations are related by one of the standard moves. Software packages such as SnapPy (SnapPy) can be used to check whether these moves are valid and can also list valid moves. We can hence represent the workflow as follows:
@@
  • To show two objects are inequivalent, one defines invariants, which are typically algebraic objects (numbers, polynomials, or groups) that are unchanged by the above moves. These moves often also capture important topological properties of these objects. The most classical invariants are homology groups and the fundamental group. The Alexander polynomial is a knot invariant derived from the fundamental group of the knot complement. Representations of the knot group give rise to the more modern twisted Alexander polynomials. A recent knot invariant rooted in representation theory is Khovanov homology (Khovanov:2000categorification). Floer homology (OzsvathSzabo:2004HF) and gauge theory (Witten:1994monopoles) give rise to highly sophisticated knot and 3- and 4-manifold invariants. These invariants often give lower bounds on hard-to-compute topological quantities such as the 3- or 4-genus of a knot. SnapPy is capable of computing many of these knot invariants when run in SageMath (sagemath). A typical workflow would thus be as follows: